OpenStudy (fibonaccichick666):

So, I have my basic analysis final today, and quite honestly, I need help. So for my first question, can someone help me make an outline for an epsilon delta proof?

OpenStudy (fibonaccichick666):

ok so this is what I have: \(set~\epsilon>0\) \(Define~ \delta :=n\epsilon~~~~ (have~to ~include~max~here~somehow?)~~s.t.~~~|x-c|<\delta \) Then \[|f(x)-L| \le\]...\[~~~~~~~|x-c|\le \epsilon \]

OpenStudy (fibonaccichick666):

@zzr0ck3r , Does that work? And the final should be epsilon/delta

OpenStudy (anonymous):

tsk

OpenStudy (anonymous):

if only dan815 were here to save us

OpenStudy (anonymous):

the way to make the perfect user name is take your hand spread it open wide and scramble it across the keyboard

OpenStudy (fibonaccichick666):

the ... is algebra to get to the point where we have delta

OpenStudy (zzr0ck3r):

\(\forall x\in D \ \forall \epsilon >0 \ \exists\delta >0 \ \forall x\in D \ (|x-a|<\delta \implies|f(x)-f(a)|<\epsilon )\)

OpenStudy (zzr0ck3r):

This is continuity

OpenStudy (fibonaccichick666):

but is that ok for just limit?

OpenStudy (zzr0ck3r):

ahh well we need a few quantifiers

OpenStudy (fibonaccichick666):

ok, so should I use set or for every for epsilon, or does it matter?

OpenStudy (zzr0ck3r):

\( \forall \epsilon >0 \ \exists\delta >0 \ \forall x\in D-\{c\} \ (|x-c|<\delta \implies|f(x)-L|<\epsilon ) \)

OpenStudy (fibonaccichick666):

what is the x an element of D for?

OpenStudy (fibonaccichick666):

ohoh oh nvm

OpenStudy (zzr0ck3r):

the -{c} might not be in the definition you use

OpenStudy (zzr0ck3r):

It does not matter

OpenStudy (fibonaccichick666):

ok so where does the delta being in the max{} belong?

OpenStudy (zzr0ck3r):

well normally we use min not max

OpenStudy (zzr0ck3r):

since often we need to have multiple conditions on |x-c|

OpenStudy (fibonaccichick666):

oh oops...

OpenStudy (zzr0ck3r):

you have it backwards, normally its like this \(|x-c|\implies ....\) Since delta is less than blah \(.......\) thus \(|f(x)-L|< \epsilon\)

OpenStudy (fibonaccichick666):

I guess I never understood how to write the bounds. Like if it is from {1, epsilon/2},

OpenStudy (zzr0ck3r):

Say we want to prove that the limit of f(x) = 3x as x goes to 3

OpenStudy (fibonaccichick666):

ok so first we guess L and do the side work to find delta right?

OpenStudy (zzr0ck3r):

right, L = 9

OpenStudy (fibonaccichick666):

so then \[|3x-9| \le 3|x-3|\]\[~~~~~~~~~~~~\le \epsilon/3\]

OpenStudy (fibonaccichick666):

so we define delta as epsilon over 3

OpenStudy (fibonaccichick666):

then we write up the whole thing.... right?

OpenStudy (zzr0ck3r):

we want| \(|3x-9|<\epsilon\) so \(|3x-3^2| = 3|x-3|\) pf: Let \(\epsilon > 0\) Choose delta such that \(\delta <\frac{\epsilon}{3}\) Then \(|x-3|<\delta \implies |x-3|<\frac{\epsilon}{3}\implies |f(x)-9|=|3x-9|=3|x-3|<\epsilon\)

OpenStudy (zzr0ck3r):

Sorry I didn't read what you wrote

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

Ok but what about something like f(x)=x^2

OpenStudy (fibonaccichick666):

then we have to use the conjugate trick

OpenStudy (zzr0ck3r):

then we will have more than one delta

OpenStudy (zzr0ck3r):

nah that's with sqrt

OpenStudy (fibonaccichick666):

ok so \[|x^2-c^2|\le |x+c||x-c|\] Then we have to bound x+c

OpenStudy (fibonaccichick666):

but on the reals, we can't make a bound... can we?

OpenStudy (zzr0ck3r):

\(|x^2-3^2|=|x-3||x+3|\) So we will do a few things Suppose \(|x-3|<1\) then \(x-3<1\implies x+3<7\)

OpenStudy (zzr0ck3r):

yes

OpenStudy (fibonaccichick666):

ok so wlog we can assume that |x+c|<1

OpenStudy (zzr0ck3r):

now we have choose \(\delta , \min(\frac{\epsilon}{7}, 1)\)

OpenStudy (zzr0ck3r):

\(\delta < \min(\frac{\epsilon}{7}, 1) \)

OpenStudy (fibonaccichick666):

how did you choose the 7?

OpenStudy (zzr0ck3r):

I wanted |x+3| < something

OpenStudy (zzr0ck3r):

the choice of 1, led to the 7

OpenStudy (fibonaccichick666):

what if you had a general case?

OpenStudy (zzr0ck3r):

?

OpenStudy (fibonaccichick666):

like instead of 3, you had c?

OpenStudy (zzr0ck3r):

and we want to get get \(|x+c|<\)something?

OpenStudy (fibonaccichick666):

right

OpenStudy (zzr0ck3r):

\(|x-c|<1\implies x-c<1\implies x+c<1+2c\implies |x+c|<1+2c \)

OpenStudy (fibonaccichick666):

so then we would use 1+2c?

OpenStudy (fibonaccichick666):

well epsilon over 1+2c?

OpenStudy (zzr0ck3r):

well, then you got to remember, that you will be replacing |x+3| with that in the string of inequalities, so you will have, most likely, something like \(\delta < \min(1, \frac{\epsilon}{1+2c})\)

OpenStudy (zzr0ck3r):

no, delta < 1 gives us |x+3|< 1+2c

OpenStudy (zzr0ck3r):

errrrr

OpenStudy (zzr0ck3r):

replace the 3's with c's lol

OpenStudy (fibonaccichick666):

ok, so for the general, we get 1+2C because we bounded |x-c|, yes?

OpenStudy (zzr0ck3r):

yes

OpenStudy (fibonaccichick666):

ok, and the min is placed directly after we define delta in the prffof?

OpenStudy (fibonaccichick666):

oh goodness, proof**

OpenStudy (fibonaccichick666):

lol

OpenStudy (zzr0ck3r):

we define delta with min

OpenStudy (fibonaccichick666):

ohhhh ok

OpenStudy (zzr0ck3r):

in most cases....

OpenStudy (zzr0ck3r):

you could equaly choose deta = min{}

OpenStudy (zzr0ck3r):

not <....

OpenStudy (fibonaccichick666):

... ya know, conceptually I get this stuff, but for some reason, this one proof, has never sunk in

OpenStudy (fibonaccichick666):

we have to define delta>0 first though, right?

OpenStudy (zzr0ck3r):

yeah we should say \(0< \delta < \min \{....\}\)

OpenStudy (zzr0ck3r):

we can say choose a delta in that interval, or use = instead of < on the rhs

OpenStudy (zzr0ck3r):

the stuff we pick inside the min must always be positive...

OpenStudy (zzr0ck3r):

What else you got? its 445 am ;)

OpenStudy (zzr0ck3r):

no one should be doing epsilon delta proofs in this hour

OpenStudy (fibonaccichick666):

ok and lol it's 745 here :)

OpenStudy (fibonaccichick666):

And, that settles ep delta question for limits, the only difference for continuity os we use f(c) and then I'll open a new question for my second query

OpenStudy (fibonaccichick666):

thanks :) if you have to go, I understand :)

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

god I love the west coast

OpenStudy (fibonaccichick666):

haha east COAAASSSSSSTTT is where it's at bro

OpenStudy (fibonaccichick666):

but unfortunately crappy teachers also call this place home o.o

OpenStudy (zzr0ck3r):

lol 2pac 4 life son

OpenStudy (zzr0ck3r):

jk ;P

OpenStudy (fibonaccichick666):

lol